Prove that if a complex number z is an nth root of a real number a , then so
is z .
1
Expert's answer
2020-05-19T13:46:54-0400

We can take any complex number in the form of "a+ib" ,

 take nth root in the form of "x+iy" . Here, n is any positive integer.

so, "\lparen a+ib \rparen ^{n}=x+iy"

As for De Moivre's Theorem, complex number

"z=cos\phi+isin\phi" ,

"z^{1/n}=r^{1/n}(cos\alpha+isin\alpha)" (I)

where "\alpha=(\phi+2\pi k)/n" ; "k=1,2,3,...,(n-1)"

so, we can get the  nth root of the complex number as the form of z bar:

"z=r(cos\phi+isin\phi)"

To take a general equation as

"\bar{z}=a-ib"

Polar coordinate form of this is "\bar{z}=r(cos\phi+isin\phi)"

Using De Moivre's theorem again, we get roots in the conjugate case

"\bar{z}^{1/n}=r^{1/n}(cos\alpha-isin\alpha)" (II)

From equations (I) and (II), we can get that real parts of the nth part of the equation

"Re(\bar{z}^{1/n})=Re(z^{1/n})" are the same.

Then proved.